OpenStudy (anonymous):

What are the roots of 2x+5/x=2 express in a+bi form

OpenStudy (turingtest):

multiply all by x: 2x^2+5=2x set equal to zero: 2x^2-2x+5=0 solve using the quadratic formula using a=2, b=-2, c=5: \[x=(-b \pm \sqrt{b^2-4ac})/2a=(2 \pm \sqrt{2^2-4(2)(5)})/2(2)\]\[=(2 \pm \sqrt{4-40})/4=(2 \pm \sqrt{-36} )/4=(2 \pm 6i)/4\]\[=1/2 \pm (3/2)i\]